# Prediction of Spallation Induced Transmutation Rates For Long-Lived Fission Products via Proton Accelerator

### v2

@@ Abstract
Long-lived fission products represent a major challenge in nuclear waste management due to persistent radiotoxicity over very long timescales. This study focuses on six of these fission products: Se-79, Zr-93, Tc-99, Sn-126, I-127, Cs-135. This study investigates the feasibility of spallation-driven transmutation, in which a high energy proton beam strikes a heavy spallation target to generate neutrons that induce transmutation in the fission products surrounding the target. Lead and depleted uranium are identified as the principal spallation target candidates, reflecting contrasting trade offs in neutron yield, secondary reactions, and heat generation. Simulations assess nuclide specific behavior under reactor scale inventories and practical geometric constraints. Results demonstrate that technetium, iodine, and selenium are strong candidates for transmutation using this pathway, while tin shows partial resistance but benefits from thermal flux. By contrast, zirconium is neutron-transparent and inefficient to transmute, and cesium suffers from low net reduction due to competition with lighter isotopes. Cost effectiveness is highly isotope-dependent: technetium is most favorable, whereas cesium and zirconium remain expensive. These findings highlight the advantages and limitations of spallation driven systems and motivate strategies with optimized target–blanket designs.
@@ 2.1 Long-Lived Fission Products (LLFPs)
LLFPs are a category of radioactive materials characterized by exceptionally long half-lives, ranging from hundreds of thousands to millions of years. Nuclides with longer half-lives are generally not a concern as their radioactive activity is too low to contribute much to residual radiotoxicity. These nuclides are produced in nuclear reactors as byproducts of the fission process. This paper assumes that LLFPs are isolated from the remaining spent nuclear fuel and are considered in terms of their isotopic composition. The LLFP targets are constructed according to the isotopic compositions found in spent nuclear fuel[3]. Table 1 provides the half-lives for each LLFP, as well as the fraction of mass each LLFP makes up in their isotopic composition.
Table 1: Initial isotopic composition mass fraction and half life for each LLFP [3].
@@ 3.1 PHITS
PHITS (Particle and Heavy Ion Transport System), a Monte Carlo particle transport code, is a key tool utilized extensively in this study [8]. This advanced simulation software is specifically designed to model the transport of radiation across various materials with high precision. PHITS excels at capturing the complex interactions between particles, heavy ions, and matter, including detailed representations of nuclear reactions and electromagnetic processes. In this paper, PHITS is employed to analyze the generated neutron flux spectra and to find the proton penetration depth of proton beams in a target. Every PHITS run initializes and transports 100,000 protons with the following beam characteristics: 1 GeV energy, 30 mA current, in a 1 cm radius. The simulation is run in continuous energy mode and tallies a multigroup flux with the CCFE-709 group structure.
@@ 3.2 FISPACT
FISPACT is a computational tool and database system designed to simulate the behavior of materials undergoing nuclear depletion and transmutation reactions [9]. It is capable of both time-dependent and static analyses of irradiated isotopes, providing a detailed inventory of the resulting isotopes. The system is composed of three primary elements: the core software responsible for modeling depletion and transmutation, a set of nuclear data libraries, and rigorous validation and verification protocols. The nuclear data libraries are formatted in ENDF6, with users able to supplement these with additional custom libraries if needed. FISPACT processes inputs such as the particle flux spectrum, initial composition and count of target isotopes, irradiation duration, and particle type. Its outputs include data on isotopic activity, the composition of irradiated materials, radiotoxicity levels, and other essential parameters. For this work, results from PHITS simulations, including the flux-energy distribution and the quantity of irradiated atoms, are fed into FISPACT to perform a 0-dimensional transmutation rate analysis for LLFPs with the CCFE-709 group structure.
@@ 3.3 Target-Blanket Concept
The central target is 2 m in axial length, with a diameter of 10 cm chosen to ensure utilization the majority of incident protons while remaining thin enough to allow the secondary neutrons produced in spallation reactions to escape without significant absorption. A preliminary study on target width showed that an optimal target diameter was between 10 and 15 cm. The target also features a 1 cm in radius, 1 m long channel that the proton beam is incident upon. The material choice for the spallation target has a strong impact on neutron production and therefore transmutation rates, and multiple material choices were individually analyzed with the combination of PHITS and FISPACT. Further discussion of target material choice is continued in Section 4.1.
@@
Figure 1: Geometry of tank layout for tallying neutron spectrum at different pin layers from a top down view. Light blue represents the spallation target, white represents the beam, brown represents heavy water, and dark blue is the reflector. Note that the LLFP pins surround the spallation target.
@@ 4.2 Spectral Sensitivity Analysis
Preliminary runs without the LLFP pins and purely D2O moderator show that there is no true fast region, as the neutrons are thermalized very quickly by the D2O. The region closest to the target (lead in this case) had a neutron flux that was 60% thermal (below 4 eV), and 10% fast (greater than 1 MeV). At 11 cm from the target, the flux was 78% thermal and only 3% fast.
@@
For some of the nuclides, such as Se-79, Tc-99, and I-129, increased thermalization from being further from the source has a minor impact that mostly washes out by 5 years of irradiation. For others, such as Sn-126 and Cs-135, there is clear and significant increases in transmutation amounts due to a more thermal spectrum, in the range of 15-20% increases in transmutation after 5 years. Zr-93 has more unclear response to the change in position, likely due to the general neutron-transparency that zirconium is known to exhibit [11], along with some stochastic noise from the PHITS simulation.
@@ 4.3 Combined Geometry
To this point, this analysis is missing a large component; the uranium target has a non-negligible fission rate that both produces neutrons and LLFP nuclides. If the uranium target’s production of LLFPs outweighs the increased transmutation from more spallation neutrons, then changing targets would be inadvisable. To estimate the LLFP generation in the depleted uranium target, PHITS-generated proton and neutron flux in the target was used as an input for two FISPACT runs. The neutron run considered fission due to neutrons in the target, and the proton run used TENDL incident proton data for fissions [12], as well as HEIR high-energy cascade data for fission and evaporation modeling [13].
@@ 5 Discussion
This work provides a systematic assessment that couples spallation target comparisons with spectral sensitivity based LLFP ordering within a mass constrained geometry representative of reactor scale waste inventories. The results highlight clear trade-offs between neutron yield, target material side-effects, isotope placement, and achievable burn rates. This framework provides a more realistic basis for assessing transmutation strategies and can guide future research in three directions: (i) experimental validation of optimized target–blanket designs, (ii) refinement of economic models including reprocessing and capital costs, and (iii) exploration of hybrid systems where nuclides such as Cs-135 are treated separately from the combined geometry. By integrating these aspects, accelerator-driven transmutation could evolve into a technically and economically credible approach for reducing the long-term radiotoxic burden of nuclear waste.

### v3

@@ Abstract
Long-lived fission products represent a major challenge in nuclear waste management due to persistent radiotoxicity over very long timescales. This study focuses on six of these fission products: Se-79, Zr-93, Tc-99, Sn-126, I-127, Cs-135. This study investigates the feasibility of spallation-driven transmutation, in which a high energy proton beam strikes a heavy spallation target to generate neutrons that induce transmutation in the fission products surrounding the target. Lead and depleted uranium are identified as the principal spallation target candidates, reflecting contrasting trade offs in neutron yield, secondary reactions, and heat generation. Simulations assess nuclide specific behavior under reactor scale inventories and practical geometric constraints. Results demonstrate that technetium, iodine, and selenium are strong candidates for transmutation using this pathway, while tin shows partial resistance but benefits from thermal flux. By contrast, zirconium is inefficient to transmute, and cesium suffers from low net reduction due to competition with lighter isotopes. Cost effectiveness is highly isotope-dependent: technetium is most favorable, whereas cesium and zirconium remain expensive. These findings highlight the advantages and limitations of spallation driven systems and motivate strategies with optimized target–blanket designs.
@@ 2.1 Long-Lived Fission Products (LLFPs)
LLFPs are a category of radioactive materials characterized by exceptionally long half-lives, ranging from hundreds of thousands to millions of years. Nuclides with longer half-lives are generally not a concern as their radioactive activity is too low to contribute much to residual radiotoxicity. These nuclides are produced in nuclear reactors as byproducts of the fission process. The LLFP targets are constructed according to the isotopic compositions found in spent nuclear fuel, where the spent nuclear fuel is assumed to be elementally separated[3]. The mass fractions for every isotope are listed in Supplementary Tables S1-S5, where the mass fraction of each LLFP isotope refers to its relative abundance within the isotopic composition of its parent element. Table 1 provides the half-lives for each LLFP, as well as the fraction of mass each LLFP makes up in their isotopic composition.
Table 1: Initial isotopic composition mass fraction and half life for each LLFP. The mass fraction of each LLFP isotope refers to its relative abundance within the isotopic composition of its parent element (e.g., the fraction of Se-79 among all selenium isotopes created in the nuclear spent fuel) [3].
@@ 2.2 Spallation
Recent research has explored accelerator driven and reactor based systems for the transmutation of LLFPs. Kang et al. demonstrated that coupling a 1 GeV proton beam with a compact subcritical assembly enables efficient Tc-99 transmutation. The neutron spectrum is adjusted with Be or Pb–Bi moderators, achieving high transmutation rates while focusing on trade offs between neutron yield, target activation, and heat load [8]. Chiba et al. demonstrated that a YD2 moderated fast reactor can effectively achieve net LLFP reduction in six major isotopes (Se-79, Zr-93, Tc-99, Pd-107, I-129, Cs-135) without isotope separation, showing that moderation in peripheral regions increases neutron capture probabilities and enhances transmutation efficiency [9]. Sun et al. proposed a photonuclear accelerator driven system combining Pb–Bi and Be layers for simultaneous transmutation of multiple LLFPs, finding that neutron spectra can be tuned to balance capture efficiency across isotopes, though at the cost of extremely high beam intensities [10]. Complementary work by Jin et al. examined proton, deuteron, and $\alpha$ induced spallation on thick LLFP targets, identifying energy windows (50–200 MeV) and projectile choices that maximize reaction yield while minimizing secondary product buildup, underscoring the importance of optimizing target composition and geometry for realistic implementation [11].
Although these studies demonstrate the feasibility of LLFP transmutation using accelerator driven or hybrid systems, most have been limited to geometries that do not capture the coupled effects of neutron spectrum, target composition, and spatial isotope arrangement. The present work expands on these efforts by analyzing the LLFPs within a unified spallation-driven framework. Using Monte Carlo simulations of Pb and depleted uranium spallation targets, this study analyzes the relationships between neutron yield, spectral characteristics, and isotope positioning to determine their combined impact on transmutation efficiency.
@@ 3.1 PHITS
PHITS (Particle and Heavy Ion Transport System), a Monte Carlo particle transport code, is a key tool utilized extensively in this study [12]. This advanced simulation software is specifically designed to model the transport of radiation across various materials with high precision. PHITS excels at capturing the complex interactions between particles, heavy ions, and matter, including detailed representations of nuclear reactions and electromagnetic processes. In this paper, PHITS is employed to analyze the generated neutron flux spectra and to find the proton penetration depth of proton beams in a target. Every PHITS run initializes and transports 100,000 protons with the following beam characteristics: 1 GeV energy, 30 mA current, in a 1 cm radius. The simulation is run in continuous energy mode and tallies a multigroup flux with the CCFE-709 group structure. The PHITS analysis in this work uses INCL4.6 [13] and GEM [14] for high energy proton reaction data, as well as JENDL-4.0 for neutron reaction data [15].
Although substantial experimental data exist for proton induced spallation reactions in materials such as lead and tungsten [4], allowing reliable neutron yield benchmarking, the neutron induced cross sections of the LLFPs present a significant source of uncertainty. Experimental validation of these cross sections is inherently difficult, as it requires isolating small quantities of pure LLFP material from spent nuclear fuel, which is both radiologically hazardous and logistically complex. Consequently, there is a scarcity of experimental data, particularly in the fast and epithermal neutron energy ranges. For example, the neutron capture cross section of Sn-126 has only two experimental data points, leaving the evaluated libraries (such as TENDL, JENDL, and ENDF) to rely almost entirely on theoretical model predictions [16] [17]. Some work exists on quantifying proton reaction uncertainties via cross-section perturbation, but only in the energy ranges of 8-30 MeV [18]. This lack of experimental validation limits the accuracy of predicted transmutation rates and underscores the need for improved nuclear data to support future design and optimization of spallation-driven transmutation systems.
@@ 3.2 FISPACT
FISPACT is a computational tool and database system designed to simulate the behavior of materials undergoing nuclear depletion and transmutation reactions [19]. It is capable of both time-dependent and static analyses of irradiated isotopes, providing a detailed inventory of the resulting isotopes. The system is composed of three primary elements: the core software responsible for modeling depletion and transmutation, a set of nuclear data libraries, and rigorous validation and verification protocols. The nuclear data libraries are formatted in ENDF6, with users able to supplement these with additional custom libraries if needed. FISPACT processes inputs such as the particle flux spectrum, initial composition and count of target isotopes, irradiation duration, and particle type. Its outputs include data on isotopic activity, the composition of irradiated materials, radiotoxicity levels, and other essential parameters. For this work, results from PHITS simulations, including the flux-energy distribution and the quantity of irradiated atoms, are fed into FISPACT to perform a 0-dimensional transmutation rate analysis for LLFPs with the CCFE-709 group structure. FISPACT simulations used TENDL-2017 data for neutron reactions [20], as well as HEIR-01 data for proton reactions [21].
@@ 3.3 Target-Blanket Concept
The central target is 2 m in axial length, with a diameter of 10 cm chosen to ensure utilization the majority of incident protons while remaining thin enough to allow the secondary neutrons produced in spallation reactions to escape without significant absorption. A preliminary study on target width showed that an optimal target diameter was between 10 and 15 cm. The target also features a 1 cm in radius, 1 m long channel that the proton beam is incident upon. The point of the void region is for the beam to interact with the spallation target in the middle of the tank instead of on the outer surface of the spallation target. The beam travels through the void region, and interacts with the spallation target, leading to the previously mentioned desired neutron production. The material choice for the spallation target has a strong impact on neutron production and therefore transmutation rates, and multiple material choices were individually analyzed with the combination of PHITS and FISPACT. Further discussion of target material choice is continued in Section 4.1.
@@
Figure 1: Geometry of tank layout for tallying neutron spectrum at different pin layers from a top down view (left) and a side view (right). Light blue represents the spallation target, white represents the beam, brown represents heavy water, and dark blue is the reflector. Note that the LLFP pins surround the spallation target and that the beam travels from left to right on the z-axis.
@@ 4.2 Spectral Sensitivity Analysis
Preliminary runs without the LLFP pins and purely D2O moderator show that there is no true fast region, as the neutrons are thermalized very quickly by the D2O. The region closest to the target (lead in this case) had a neutron flux that was 60% thermal (below 4 eV), and 10% fast (greater than 1 MeV). At 11 cm from the target, the flux was 78% thermal and only 3% fast. A figure showing the neutron flux spectrum at different radial positions can be found in the supplementary information, Figure S1.
@@
For some of the nuclides, such as Se-79, Tc-99, and I-129, increased thermalization from being further from the source has a minor impact that mostly washes out by 5 years of irradiation. For others, such as Sn-126 and Cs-135, there is clear and significant increases in transmutation amounts due to a more thermal spectrum, in the range of 15-20% increases in transmutation after 5 years. Zr-93 has more unclear response to the change in position, likely because the scattering cross-section dominates the total cross-section, along with some stochastic noise from the PHITS simulation.
@@ 4.3 Combined Geometry
Analysis of the impact of the irradiation with different target materials and ordering on radiotoxicity of the LLFP materials can be found in the Supplementary materials, specifically Figures S2, S3, and Table S6. Radiotoxicity here is defined as the ingestion dose retrieved from the FISPACT output, which corresponds to ICRP dose coefficients [23].
@@
To this point, this analysis is missing a large component; the uranium target has a non-negligible fission rate that both produces neutrons and LLFP nuclides. If the uranium target’s production of LLFPs outweighs the increased transmutation from more spallation neutrons, then changing targets would be inadvisable. To estimate the LLFP generation in the depleted uranium target, PHITS-generated proton and neutron flux in the target was used as an input for two FISPACT runs. The neutron run considered fission due to neutrons in the target, and the proton run used TENDL incident proton data for fissions, as well as HEIR high-energy cascade data for fission and evaporation modeling.
@@ 5 Discussion
Some limitations of this work are as follows: the spallation sources is assumed constant for 5 years for all of the FISPACT runs, and cross-section data does not have uncertainty included in the results, something that should be clarified in future work. There is also a lack of experimental neutron cross section data for the LLFPs. This work provides a systematic assessment that couples spallation target comparisons with spectral sensitivity based LLFP ordering within a mass constrained geometry representative of reactor scale waste inventories. The results highlight clear trade-offs between neutron yield, target material side-effects, isotope placement, and achievable burn rates. This framework provides a more realistic basis for assessing transmutation strategies and can guide future research in three directions: (i) experimental validation of optimized target–blanket designs, (ii) refinement of economic models including reprocessing and capital costs, and (iii) exploration of hybrid systems where nuclides such as Cs-135 are treated separately from the combined geometry. By integrating these aspects, accelerator-driven transmutation could evolve into a technically and economically credible approach for reducing the long-term radiotoxic burden of nuclear waste.